(Appendix 12A) Marvel Company estimates that the following costs and activity would be associated with the manufacture and sale of one unit of product Y:

Number of Units Sold Annually 20,000
Required Investment $400,000
Unit Product Cost $25
Selling, General, and Administrative Expenses $130,000



If the company uses the absorption costing approach to cost-plus pricing and desires a 15% rate of return on investment (ROI), what would be the required markup on absorption cost for product Y?

Multiple Choice

  • 12%.

  • 15%.

  • 26%.

  • 38%.
Answer #1

Profit expected = 400,000*15% = 60,000

Gross margin expected = 60,000 + 130,000 = 190,000

Gross margin expected per unit = 190,000/20,000 = 9.5

Markup = 9.5/25

= 38%

Option D is the answer